# MIVE: New Design and Benchmark for Multi-Instance Video Editing

### v1

@@ 4.1 MIVE Dataset Construction
Available datasets are unsuitable for multi-instance video editing (VE) tasks. TGVE [59] and TGVE+ [47] contain a limited number of videos, have few objects with few instances per object class, and have no instance masks. The datasets (subsets of DAVIS [39], TGVE [59], or Internet videos) used in GAV [25] and EVA [63] are either partially accessible or unavailable. DAVIS [39] treats multiple instances of the same object class as a single entity and is, thus, not suitable for multi-instance VE.
@@
Table 1 compares our MIVE Dataset to other VE datasets. EVA’s dataset is not publicly available, so we use the statistics in their paper. Compared to others, MIVE Dataset offers the most number of evaluation videos and the greatest diversity in terms of the number of objects per video and the number of instances per object class. Our dataset also demonstrates a wide range of instance sizes, from small objects covering as little as 25 pixels ($\sim$0.01%) per video to large objects occupying $\sim$98.68of the video. See Supp. Sec. C.1 for more details.
@@ 4.2 Cross-Instance Accuracy Score
SpaText [1] and InstanceDiffusion [55] assess local textual alignment using the Local Textual Faithfulness, the cosine similarity between CLIP [41] text embeddings of an instance caption and image embeddings of the cropped instance. While this measures text-instance alignment, it overlooks potential cross-instance information leakage, where a caption for an instance influences others. We observe that this score is sometimes higher for instances that should be unaffected by an instance caption (see Supp. Sec. C.2).
@@ 5 Experiments
Scope GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ BP $\downarrow$ TC $\uparrow$ TF $\uparrow$ Leakage $\uparrow$
ControlVideo [70] ICLR’24 Global 0.9743 0.2738 0.8856 0.9548 0.1960 0.4941 0.4967 72.8690 6.97 14.00 6.26
FLATTEN [11] ICLR’24 Global 0.9679 0.2388 0.2637 0.9507 0.1881 0.2469 0.5111 62.8136 32.32 2.45 9.74
RAVE [26] CVPR’24 Global 0.9675 0.2727 0.5777 0.9551 0.1869 0.3512 0.4945 64.8703 10.45 3.61 4.52
TokenFlow [16] ICLR’24 Global 0.9686 0.2569 0.5622 0.9478 0.1868 0.3501 0.5307 68.6688 7.61 3.16 4.26
FreSCo [62] CVPR’24 Global 0.9541 0.2527 0.4202 0.9324 0.1860 0.2962 0.5172 85.1843 3.55 1.81 3.42
GAV [25] ICLR’24 Local, Multiple 0.9660 0.2566 0.5504 0.9514 0.1893 0.3703 0.5492 60.0773 8.90 7.74 9.74
MIVE (Ours) - Local, Multiple 0.9604 0.2750 0.8557 0.9478 0.2138 0.6419 0.7100 54.3452 30.20 67.23 62.06
@@
Implementation details. We conduct our experiments on a single NVIDIA RTX A6000 GPU. We use Stable Diffusion [44] v1.5 with ControlNet [69] and depth [43] as condition. We perform $T=50$ DDIM denoising steps after doing 100 inversion steps with an empty text following [11]. Our CFG [19] scale is 12.5. We use RAFT [51] to estimate the optical flows for FLATTEN [11] that we adopt in our framework. We apply our IPR and alternate between LPS and NPS for the first 40 steps, switching to only NPS for the last 10 steps. Specifically, in the first 40 steps, we apply 9 LPS steps and and 1 NPS step with $l=3$ re-inversion steps, repeating this cycle four times. In the last 10 steps, we deactivate IPR and perform NPS with $l=2$ re-inversion steps.
@@
Global metrics assess overall frame quality but they overlook nuances in individual instance edits, essential for multi-instance tasks. To address this, we use Local Temporal Consistency (LTC), Local Textual Faithfulness (LTF) [55, 1], and Instance Accuracy (IA) computed using the cropped edited instances. Metric computation details are in Supp. Sec. C.3. We also quantify cross-instance leakage through our proposed CIA (Sec. 4.2) and background leakage through Background Preservation (BP) [71] that is the $L_{1}$ distance between input and edited frame backgrounds.
@@ 5.1 Experimental Results
Figure 5: Qualitative comparison for three videos (with increasing difficulty from left to right) in our MIVE dataset. (a) shows the color-coded masks overlaid on the input frames to match the corresponding instance captions. (b)-(f) use global target captions for editing. (g) uses global and instance target captions along with bounding boxes (omitted in (a) for better visualization). Our MIVE in (h) uses instance captions and masks. Unfaithful editing examples are shown in red arrow and attention leakage are shown in green arrow.
@@
We compare our framework against recent zero-shot video editing (VE) methods: (i) five global editing: ControlVideo [70], FLATTEN [11], RAVE [26], FreSCo [62], TokenFlow [16], and (ii) one multi-object editing: GAV [25]. We use single global source and edit captions for the global editing methods. We use a combination of global and local source and target captions along with bounding box conditions for GAV. We exclude the concurrent work EVA [63] as the code is not yet publicly available and Video-P2P [32] since it edits one instance at a time, accumulating error when used in multi-instance editing scenarios (see Supp. Sec. D.4). For each baseline, we use their default settings across all videos.
@@
Qualitative comparison. We present qualitative comparisons in Fig. 5. As shown, ControlVideo and GAV suffer from attention leakage (e.g. “yellow” affecting the “alien” in Video 1 of Fig. 5-(b) and -(g)) while FLATTEN, RAVE, TokenFlow, FreSCo, and GAV exhibit unfaithful editing for all the examples (Fig. 5-(c) to Fig. 5-(g)). Moreover, ControlVideo and FreSCo exhibit a mismatch in editing instances as shown in Fig. 5-(b), Video 2 (incorrectly turning the man on the left to “woman in a red dress”) and Fig. 5-(f), Video 1 (incorrectly turning the washing machine on the left to “yellow washing machine”), respectively. In contrast, our method confines edits within the instance masks, faithfully adheres to instance captions, and does not exhibit mismatch in instance editing.
Quantitative comparison. Table 2 shows the quantitative comparisons. The SOTA methods yield (i) low LTF and IA scores, indicating less accurate instance-level editing, and (ii) low CIA scores, reflecting significant attention leakage. However, they score better in temporal consistency-related metrics (GTC and LTC) due to either (i) smooth results with less textures and details (Fig. 5-(b), Video 2) or (ii) unfaithful editing (Fig. 5-(c) to Fig. 5-(g)) that simply reconstructs the input video, preserving structure and motion without altering the instances. In contrast, our MIVE achieves top scores in GTF and key multi-instance VE metrics (LTF, IA, CIA, and BP), while maintaining a competitive LTC score. Our FA ranks second to ControlVideo, which we attribute to CLIP’s limitations in compositional reasoning [23, 36], preventing it from capturing complex relationships between the instances in the global caption and frames. This potentially leads to higher FA scores despite mismatches in the instance edits (e.g., edits are switched for the two people in Video 2 of Fig. 5-(b)). Note that our MIVE also shows the best BP performance without using an inpainting model, as GAV does, indicating its effectiveness in isolating instance edits while preserving and preventing leakage in the background. Additional demo videos are available in the Supp.
Methods LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA $\uparrow$ BP $\downarrow$
DMS Only NPS 0.9460 0.2072 0.5587 0.6663 54.6597
Only LPS 0.9483 0.2068 0.5716 0.6688 50.8549
LPS + NPS w/o Re-Inversion 0.9485 0.2080 0.5776 0.6783 52.3240
Ours, Full 0.9478 0.2138 0.6419 0.7100 54.3452
IPR No Modulation [44] 0.9535 0.2060 0.5225 0.6553 50.1319
Dense Diffusion [28] 0.9482 0.2136 0.6215 0.6891 59.2100
Ours, Full 0.9478 0.2138 0.6419 0.7100 54.3452
@@
Table 3: Ablation study results on DMS (Sec. 3.2) and IPR (Sec. 3.3). LPS and NPS denotes Latent Parallel Sampling and Noise Parallel Sampling, respectively. The best and second best scores are shown in red and blue, respectively.
@@
User study. We conduct a user study to compare our method with baselines and report the results in Tab. 2. We select 50 videos from our dataset covering diverse scenarios (varying number of objects per clip, number of instances per object class, and instance size). We task 31 participants to select the method with the best temporal consistency (TC), textual faithfulness (TF), and minimal editing leakage (Leakage). Our method outperforms baselines, with a 67.23% chance of being selected as the best for TF and 62.06% for minimal leakage. MIVE also achieves second best TC, with a 30.20% selection rate. This discrepancy with our GTC and LTC scores may stem from CLIP’s focus on consecutive frame consistency, while human evaluators prioritize the overall visual stability across instances and frames. Details of our user study are in the Supp.
@@ 5.2 Ablation Studies
For our Ablation Studies, we report key multi-instance VE metrics in Tab. 3 and provide the Global Scores in the Supp.
Ablation study on DMS. We conduct an ablation study on our DMS, presenting qualitative results in Fig. 6 and quantitative results in Tab. 3. From Fig. 6, we observe the following: (b) using only NPS results in suboptimal and unfatifhul editing; (c) using only LPS produces noisy, blurry textures with poor details; (d) performing alternating LPS and NPS produces similar artifacts, but slightly sharper edges compared to (c), emphasizing the importance of alternating LPS and NPS; and (e) adding re-inversion produces faithful edits with less artifacts, sharper boundaries, and enhanced details, demonstrating the effectiveness of our re-inversion technique. Table 3 reinforces these findings, with our full method achieving the highest editing faithfulness, as reflected in LTF, IA, and CIA, though with a slight reduction in BP and LTC. Overall, our final model strikes a balanced trade-off between faithfulness, BP and LTC.
Figure 7: Ablation study on IPR (Sec. 3.3).
Ablation study on IPR. We also conduct an ablation study on our IPR, with qualitative results presented in Fig. 7 and quantitative results in Tab. 3. As shown in Fig. 7-(b), the model tends to reconstruct the input frames when modulation is omitted. The reconstruction thereby preserves the high LTC of the input but greatly suffers from unfaithful editing (low LTF and IA) and leakage (low CIA). Introducing modulation via DenseDiffusion [28] (Fig. 7-(c)) slightly reduces LTC but enhances faithfulness and produces less leakage, though it produces severe blurring and smoothing artifacts due to the change in spatial distribution of attention maps. These artifacts in turn produce worse BP. In contrast, Our IPR produces faithful edits (best LTF and IA) and less leakage (best CIA), while also producing less artifacts (better BP) than DenseDiffusion with slight reduction in LTC.
@@ 6 Conclusion
In this work, we introduce MIVE, a novel general-purpose mask-based multi-instance video editing framework featuring Disentangled Multi-instance Sampling (DMS) and Instance-centric Probability Redistribution (IPR). Our method achieves faithful and disentangled edits with minimal attention leakage, outperforming existing SOTA methods in both qualitative and quantitative analyses on our newly proposed MIVE Dataset. We further propose a new Cross-Instance Accuracy (CIA) Score to quantify attention leakage. Our user study supports MIVE’s robustness and effectiveness with participants favoring our method.
@@ A Preliminaries
B Demo Videos
We provide sample input and edited videos for our method along with baseline methods in our project page: https://kaist-viclab.github.io/mive-site/.
C Dataset and Metrics Additional Details
@@
For User Study 50 $13-46$ 2-9 38 1-16 $0.05\sim 75.62$
For Demo 40 $13-46$ 2-9 35 1-16 $0.05\sim 69.74$
Table 4: Statistics of our multi-instance video editing dataset in various use cases. (i) MIVE Dataset is the full set of our dataset including both “stuff” and “thing” categories. (ii) For Editing, we only edit objects in the “thing” categories, thus, decreasing some statistics. (iii) For User Study, we only select 50 videos that cover diverse scenarios. (iv) For the Demo, we only select 40 videos that cover diverse scenarios.
@@
C.1 MIVE Dataset Construction
@@
C.2 Cross-Instance Accuracy (CIA) Score
@@
Our Cross-Instance Accuracy (CIA) Score is proposed to address the shortcomings of existing video editing metrics, particularly their inability of accounting for potential editing leakage (commonly termed as attention leakage in literatures). The Global Textual Faithfulness (GTF) and Frame Accuracy (FA) metrics cannot account for cases where the text prompt for one instance leaks to another instance, as they cannot capture the nuances of individual instances within the global captions and frames. The Instance Accuracy (IA) only determines whether a cropped instance is more aligned with its target caption compared to its source caption, without considering whether or not another instance caption is affecting the cropped instance. The Local Textual Faithfulness (LTF) only quantifies the alignment of the instance’s target caption and its corresponding cropped instance, but it also overlooks potential attention leakage from other instance captions. While Background Preservation (BP) can measure modifications and attention leakage in the background, it does not take into consideration the leakage in instances that should be unaffected by a given instance caption. We further observe that the Local Textual Faithfulness between a cropped instance and another target instance caption, which should not affect the cropped instance, is sometimes higher than the score between the cropped instance and its corresponding target instance caption (red text in Fig. 11).
@@
C.3 Local Metrics Computation

### v2

@@ 4.1 MIVE Dataset Construction
Existing datasets are unsuitable for multi-instance video editing (VE) tasks. TGVE [60] and TGVE+ [49] contain a limited number of videos, feature few objects with few instances per object class, and lack instance masks. Similarly, the dataset used in VideoGrain [65] includes only a few objects with limited instances per class. GAV’s [28] dataset is only partially available, and DAVIS [41] does not distinguish between multiple instances of the same object class, making it unsuitable for multi-instance VE.
@@
Table 1 compares our MIVE Dataset to other VE datasets. Compared to others, MIVE Dataset offers the most number of evaluation videos and the greatest diversity in terms of the number of objects per video and the number of instances per object class. Our dataset also demonstrates a wide range of instance sizes, from small objects covering as little as 25 pixels ($\sim$0.01%) per video to large objects occupying $\sim$98.68of the video. See Supp. for more details.
@@ 4.2 Cross-Instance Accuracy Score
SpaText [1] and InstanceDiffusion [55] assess local textual alignment using the Local Textual Faithfulness, the cosine similarity between CLIP [43] text embeddings of an instance caption and image embeddings of the cropped instance. While this measures text-instance alignment, it overlooks potential cross-instance information leakage, where a caption for an instance influences others. We observe that this score is sometimes higher for instances that should be unaffected by an instance caption (see Supp.).
@@ 5 Experiments
Method Venue Editing Local Scores Leakage Scores User Study
Scope LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ TC $\uparrow$ TF $\uparrow$ Leakage $\uparrow$
ControlVideo [72] ICLR’24 Global 0.9548 0.1960 0.4944 0.4963 0.5327 0.4242 0.0167 0.0656 0.0208
FLATTEN [14] ICLR’24 Global 0.9507 0.1881 0.2463 0.5109 0.8220 0.1314 0.3260 0.0594 0.1646
RAVE [30] CVPR’24 Global 0.9551 0.1869 0.3504 0.4950 0.7699 0.2188 0.0667 0.0354 0.0458
GAV [28] ICLR’24 Local, Multiple 0.9518 0.1895 0.3733 0.5516 0.8618 0.0914 0.0792 0.0938 0.1000
VideoGrain [65] ICLR’25 Local, Multiple 0.9423 0.2026 0.4798 0.5868 0.8929 0.0488 0.0802 0.1177 0.1542
PRIMEdit (Ours) - Local, Multiple 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586 0.3365 0.5687 0.4510
@@
Implementation details. We evaluate our PRIMEdit framework on our MIVE dataset (Sec. 4.1), editing 12-32 frames per video, and conduct our experiments on a single NVIDIA RTX A6000 GPU. We implement our PRIMEdit on top of a baseline that integrates Stable Diffusion [46] v1.5, AnimateDiff [21] motion modules and ControlNet [71] with depth [45] as our condition. We perform $T=50$ DDIM denoising steps after doing 100 inversion steps with an empty text following [14]. Our CFG [24] scale is 12.5. We apply our IPR in the first $10\%$ of the total denoising steps and switch to vanilla cross-attention [46] in the remaining steps. We perform SNS for the first $40\%$ of the denoising steps and perform PNS for the remaining steps. Our re-inversion steps $l$ is set to $l=2$.
@@
Global metrics assess overall frame quality but they overlook nuances in individual instance edits, essential for multi-instance tasks. To address this, we use Local Temporal Consistency (LTC), Local Textual Faithfulness (LTF) [55, 1], and Instance Accuracy (IA) computed using the cropped edited instances. Metric computation details are in the Supp. We also quantify cross-instance leakage through our proposed CIA (Sec. 4.2) and background leakage through SSIM [58] and LPIPS [13, 29] scores.
@@ 5.1 Experimental Results
Figure 5: Qualitative comparison for three videos (with increasing difficulty from left to right) in our MIVE dataset. (a) shows the color-coded masks overlaid on the input frames to match the corresponding instance captions. (b)-(d) use global target captions for editing. (e) uses global and instance target captions along with bounding boxes (omitted in (a) for better visualization). (f) uses masks and global and local target captions. Our PRIMEdit in (g) uses instance captions and masks. Unfaithful editing examples are shown in red arrow and attention leakage are shown in green arrow.
@@
We compare our PRIMEdit against recent zero-shot video editing (VE) methods: (i) five global editing: ControlVideo [72], FLATTEN [14], RAVE [30], FreSCo [63], TokenFlow [20], and (ii) two multi-object editing: GAV [28] and VideoGrain [65]. We show results for the first 3 global and 2 multi-object video editing baselines in this main paper and present the rest in the Supp. We use (i) single global source and target captions for the global editing methods and (ii) a combination of global and local source and target captions along with bounding box and mask conditions for GAV and VideoGrain. Our PRIMEdit requires object masks and local target captions only. We exclude Video-P2P [36] since it edits one instance at a time, accumulating error when used in multi-instance editing scenarios (see Supp.). For each baseline above, we use its default settings.
@@
Qualitative comparison. We present qualitative comparisons in Fig. 5. As shown, ControlVideo, GAV, and VideoGrain suffer from attention leakage (e.g. “yellow” affecting the “alien” in Video 1 of Fig. 5-(b), (e), and (f)) while FLATTEN, RAVE, and GAV exhibit unfaithful editing for all the examples (Fig. 5-(c) to Fig. 5-(e)). Moreover, ControlVideo exhibit a mismatch in editing instances as shown in Fig. 5-(b), Video 2 (incorrectly turning the man on the left to “woman in a red dress”). In contrast, our PRIMEdit confines edits within the instance masks, faithfully adheres to instance captions, and does not exhibit mismatch in instance editing.
Quantitative comparison. Table 2 shows the quantitative comparisons. We show only the local and leakage scores along with the user study results in this main paper and present the global scores in the Supp. The SOTA methods yield (i) low LTF and IA scores, indicating less accurate instance-level editing, and (ii) low CIA scores, reflecting significant attention leakage. In contrast, our PRIMEdit achieves the best scores in key multi-instance VE metrics (LTC, LTF, IA, and CIA) demonstrating its superior ability to maintain temporal consistency, fine-grained textual alignment, and instance-aware modifications. Note also that our PRIMEdit exhibits strong performance in background preservation metrics, achieving the best SSIM and second-best LPIPS scores without relying on an inpainting model like GAV [28]. This further underscores its effectiveness in isolating instance edits while preserving and preventing leakage in the background. We provide more qualitative comparisons in the demo videos in the Supp.
Methods LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
IPR No Modulation [46] 0.9564 0.2018 0.4920 0.6497 0.9007 0.0600
Dense Diffusion [32] 0.9530 0.2078 0.5845 0.6774 0.8977 0.0604
Ours, Full 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
DMS Only SNS 0.9503 0.2047 0.5288 0.6743 0.9063 0.0500
Only PNS 0.9542 0.2044 0.5311 0.6718 0.8993 0.0585
SNS + PNS (no re-inv) 0.9535 0.2051 0.5323 0.6728 0.9040 0.0525
Ours, Full 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
@@
Table 3: Ablation study results on IPR (Sec. 3.2) and DMS (Sec. 3.3). SNS, PNS, and re-inv denotes Series Noise Sampling, Parallel Noise Sampling, and Re-Inversion, respectively. The best and second best scores are shown in red and blue, respectively.
@@
User study. We conduct a user study to compare our PRIMEdit with the SOTA baselines, and report the results in Tab. 2. We select 30 videos from our dataset covering diverse scenarios (varying number of objects per clip, number of instances per object class, and instance size). We task 32 participants to select the method with the best temporal consistency (TC), textual faithfulness (TF), and minimal editing leakage (Leakage). Our PRIMEdit demonstrates clear superiority against previous SOTA methods with 33.65% chance of being selected as the best for TC, 56.87% for TF, and 45.10% for the least amount of leakage. Details of our user study are in the Supp.
@@ 5.2 Ablation Studies
In our ablation studies, we report key multi-instance VE metrics in Tab. 3 and provide the Global Scores in the Supp.
Figure 6: Ablation study on IPR (Sec. 3.2).
Ablation study on IPR. We conduct an ablation study on our IPR, with qualitative results presented in Fig. 6 and quantitative results in Tab. 3. As shown in Fig. 6-(b), the model tends to reconstruct the input frames when modulation is omitted. The reconstruction thereby preserves the high LTC of the input but suffers from unfaithful editing (low LTF and IA) and leakage (low CIA). Introducing cross-attention modulation via DenseDiffusion [32] reduces LTC but improves faithfulness and minimizes leakage. However, as presented in the green box in Fig. 6-(c) and [56], it struggles to generate visually appealing results when the computationally expensive spatio-temporal attention modulation is absent. In contrast, our IPR effectively reduces background artifacts (best SSIM and LPIPS), maintains temporal consistency (second-best), enhances edit faithfulness (second-best LTF and IA), and minimizes leakage (second-best CIA), all while operating solely on the cross-attention. We provide cross attention visualization in the Supp.
@@
Ablation study on DMS. We also conduct an ablation study on our DMS, presenting qualitative results in Fig. 7 and quantitative results in Tab. 3. From Fig. 7 and Tab. 3, we observe the following: (i) using only SNS prevents excessive leakage (high CIA) but suffers from temporal inconsistency (lowest LTC) as shown in Fig. 7-(b), green arrow; (ii) using only PNS improves LTC as objects are fused early in the denoising process but introduces some leakage (decrease in CIA), as seen in Fig. 7-(c), red arrow, where the green color from the plants leaks onto the man’s face; (iii) performing SNS followed by PNS mitigates the issues observed in (i) and (ii), improving both CIA and LTC; and (iv) adding re-inversion improves IA and boosts LTC by harmonizing instances with a slight decline in CIA. Overall, our final model strikes a balanced trade-off between LTC, faithfulness, leakage, and background preservation.
@@ 6 Conclusion
In this work, we introduce PRIMEdit, a novel multi-instance video editing framework featuring Instance-centric Probability Redistribution (IPR) and Disentangled Multi-instance Sampling (DMS). Our method achieves faithful and disentangled edits with minimal attention leakage, outperforming existing SOTA methods in both qualitative and quantitative analyses on our newly proposed MIVE dataset. We further propose a new Cross-Instance Accuracy (CIA) Score to quantify attention leakage. Our user study supports PRIMEdit’s robustness and effectiveness with participants favoring our method.
@@ A Preliminaries
B Dataset and Metrics Additional Details
@@
For User Study 30 $15-46$ 4-11 69 1-4 $0.05\sim 80.49$
Table 4: Statistics of our multi-instance video editing dataset in various use cases. (i) MIVE Dataset is the full set of our dataset including both “stuff” and “thing” categories. (ii) For Editing, we only edit objects in the “thing” categories, thus, decreasing some statistics. (iii) For User Study, we only select 30 videos that cover diverse scenarios.
@@
B.1 MIVE Dataset Construction
@@
B.2 Cross-Instance Accuracy (CIA) Score
@@
Our Cross-Instance Accuracy (CIA) Score is proposed to address the shortcomings of existing video editing metrics, particularly their inability of accounting for potential editing leakage (commonly termed as attention leakage in literatures). The Global Textual Faithfulness (GTF) and Frame Accuracy (FA) metrics cannot account for cases where the text prompt for one instance leaks to another instance, as they cannot capture the nuances of individual instances within the global captions and frames. The Instance Accuracy (IA) only determines whether a cropped instance is more aligned with its target caption compared to its source caption, without considering whether or not another instance caption is affecting the cropped instance. The Local Textual Faithfulness (LTF) only quantifies the alignment of the instance’s target caption and its corresponding cropped instance, but it also overlooks potential attention leakage from other instance captions. While SSIM and LPIPS can measure modifications and attention leakage in the background, they do not take into consideration the leakage in instances that should remain unaffected by a given instance caption. We further observe that the LTF between a cropped instance and another target instance caption, which should not affect the cropped instance, is sometimes higher than the score between the cropped instance and its corresponding target instance caption (red text in Fig. 11).
@@
B.3 Local Metrics Computation